what's the probability of Rolling an even number on a die and then flipping a coin that lands on heads give your answer as a reduced fraction


1
Expert's answer
2022-04-28T13:19:50-0400
"P(even)=\dfrac{3}{6}=\dfrac{1}{2}"

"P(H)=\dfrac{1}{2}"

"P(even\ and\ H)=\dfrac{1}{2}(\dfrac{1}{2})=\dfrac{1}{4}"